The sum of two consecutive integers is less than 52. Find the pair with the largest sum.
Answer by ikleyn(52798)

.
x + (x+1) = 51.

2x + 1 = 51,

2x = 51 - 1,

2x = 50,

x = 25.

The pair of two consecutive integers is (25,26), with the sum 51.